OpenStudy (anonymous):

what is the equation in slope-intercept form for the line that passes through points (-1,3) and (-2,-3)?

OpenStudy (anonymous):

find slope

OpenStudy (anonymous):

y=mx+b is the formula for slope intercept form

OpenStudy (anonymous):

(y2-y1)/(x2-x1)

OpenStudy (anonymous):

m is slope, ^^ using that you get -6

OpenStudy (anonymous):

(-3-3)/(-2--1)=-6/-1=6

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -1}}\quad ,&{\color{blue}{ 3}})\quad &({\color{red}{ -2}}\quad ,&{\color{blue}{ -3}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (anonymous):

my bad 6 lol

OpenStudy (anonymous):

y=6(x)+b 3=6(-1)+b 3=-6+b 9=b

OpenStudy (anonymous):

y=6x+9

OpenStudy (anonymous):

now plug in one of the coordinate sets to check

OpenStudy (anonymous):

Qué entiende mi bebé?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

help me please